# Training-Free Safe Denoisers for Safe Use of Diffusion Models

### v2

@@ 3.2 Practial Considerations
For computing Eq. (4), we need to compute three terms: the data denoiser $\mathbb{E}_{\text{data}}[\mathbf{x}|\mathbf{x}_{t}]$, the unsafe denoiser $\mathbb{E}_{\text{unsafe}}[\mathbf{x}|\mathbf{x}_{t}]$ and the weight $\beta^{*}(\mathbf{x}_{t})$. We approximate $\mathbb{E}_{\text{data}}[\mathbf{x}|\mathbf{x}_{t}]$ by utilizing a pre-trained diffusion model. Consequently, the task reduces to deriving $\mathbb{E}_{\text{unsafe}}[\mathbf{x}|\mathbf{x}_{t}]$ and the weight. This section delineates the approach to compute these quantities.
Unsafe denoiser Approximation.
@@
Each numerator and denominator terms of Eq. (6) approximates the numerator and denominator terms of Eq. (3), respectively. It shows that an unsafe denoiser can be expressed as a weighted sum of the unsafe dataset. Here, the weights $\{\frac{q_{t}(\mathbf{x}_{t}|\mathbf{x}^{(n)})}{\sum_{m=1}^{N}q_{t}(\mathbf{x}_{t}|\mathbf{x}^{(m)})}\}$ form a sum-to-one normalized vector across the unsafe data points, so the unsafe denoiser is approximated as a weighted unsafe data point.
Estimate of the weight.
Next, we turn our attention to the computation of $\beta^{*}(\mathbf{x}_{t})$ in Eq. (5). Direct calculation is intractable due to the denominator $Z_{\text{safe}}\int p_{\text{safe}}(\mathbf{x})q_{t}(\mathbf{x}_{t}|\mathbf{x})$, which is computationally infeasible [^1] It requies computing $q_{t}(\mathbf{x}_{t}|\mathbf{x})$ over all safe data $\mathbf{x}\sim p_{\text{safe}}(\mathbf{x})$, where safe data includes the entire training dataset excluding few unsafe data. Modern text-to-image models like Stable Diffusion (Rombach et al., 2022) are trained with billions of training data (Schuhmann et al., 2022), and is infeasible to iterate the entire data at inference time. to evaluate at every sampling steps. To address this challenge, we approximate $\beta^{*}$ as
@@
$\beta(\mathbf{x}_{t})$ $=\int p_{\text{unsafe}}(\mathbf{x})q_{t}(\mathbf{x}_{t}|\mathbf{x})\mathop{}\!\mathrm{d}\mathbf{x}$
$\approx\frac{1}{N}\sum_{n=1}^{N}q_{t}(\mathbf{x}_{t}|\mathbf{x}^{(n)})$
@@
  Input: A pre-trained diffusion model $\bm{\epsilon}_{\bm{\theta}}$; Unsafe data $\{\mathbf{x}^{(n)}\}_{n=1}^{N}$; Hyperparameters $\eta$ and $\beta_{t}$; Critical timesteps $C\subseteq[1,...,T]$; If text-conditional model, positive prompts $\mathbf{c}_{+}$ and unsafe prompts $\mathbf{c}_{US}$
  for $t=T$ to 00 do
@@
       $\mathbb{E}_{\text{data}}[\mathbf{x}|\mathbf{x}_{t},\mathbf{c}_{+}]\leftarrow\frac{1}{\alpha_{t}}\big{(}\mathbf{x}_{t}-\sigma_{t}\bm{\epsilon}_{\bm{\theta}}(\mathbf{x}_{t},t,\mathbf{c}_{+})\big{)}$
       $\mathbb{E}_{\text{data}}[\mathbf{x}|\mathbf{x}_{t},\mathbf{c}_{US}]\leftarrow\frac{1}{\alpha_{t}}\big{(}\mathbf{x}_{t}-\sigma_{t}\bm{\epsilon}_{\bm{\theta}}(\mathbf{x}_{t},t,\mathbf{c}_{US})\big{)}$
@@
       $\mathbf{x}_{0|t}\leftarrow$ Eq. (7)
@@
$\begin{split}&\mathbf{x}_{0|t}\\ &=\mathbb{E}_{\text{data}}[\mathbf{x}|\mathbf{x}_{t}]+\eta\beta(\mathbf{x}_{t})(\mathbb{E}_{\text{data}}[\mathbf{x}|\mathbf{x}_{t}]-\hat{\mathbb{E}}_{\text{unsafe}}[\mathbf{x}|\mathbf{x}_{t}]),\end{split}$ (7)
@@
3.3 Extending Safe Denoiser to Text-to-Image
Our approach can be combined with existing text-based guidance methods to enhance their performance:
$\mathbf{x}_{0|t}=\underbrace{\mathbb{E}_{\text{data}}[\mathbf{x}|\mathbf{x}_{t}]+\beta^{*}(\mathbf{x}_{t})(\mathbb{E}_{\text{data}}[\mathbf{x}|\mathbf{x}_{t}]-\mathbb{E}_{\text{unsafe}}[\mathbf{x}|\mathbf{x}_{t}]}_{\text{Safe Denoiser}})$
$\quad\quad+\underbrace{\lambda(\mathbb{E}_{\text{data}}[\mathbf{x}|\mathbf{x}_{t},\mathbf{c}_{+}]-\mathbb{E}_{\text{data}}[\mathbf{x}|\mathbf{x}_{t}]}_{\text{CFG}})$ (8)
$\quad\quad-\underbrace{\mu(\mathbf{c}_{+},\mathbf{c}_{S};\gamma,\lambda)(\mathbb{E}_{\text{data}}[\mathbf{x}|\mathbf{x}_{t},\mathbf{c}_{S}]-\mathbb{E}_{\text{data}}[\mathbf{x}|\mathbf{x}_{t}]}_{\text{SLD}}).$
Using this denoiser allows us to negate data samples based on the information from the images (from our safe denoiser) and the information based on the prompts (from both CFG and SLD). Note this Eq. (3.3) includes only the additional term for the safe denoiser compared to Eq. (2). In implementation, as described in Sec. 3.2, we approximate the second term of Eq. (3.3) by Eq. (7). In diffusion sampling, we utilize this safe $\mathbf{x}_{0|t}$ in either DDPM (Ho et al., 2020) or DDIM (Song et al., 2020), see Algorithm 1 for details.
@@ 4 Related Work
Earlier work on machine unlearning in generative modelling focused on object unlearning in classification (forgetting images from a selected class), unconditional image generation (forgetting harmful images) or concept erasing (forgetting harmful concepts). Most of the work belonging to this category required retraining the entire generative models or some part of them, rather than modifying the sampling trajectory or input prompts (Heng & Soh, 2023; Li et al., 2024; Tiwary et al., 2025; Zhang et al., 2024a; Gandikota et al., 2023b; Lu et al., 2024; Gong et al., 2024; Lu et al., 2024). In more recent work, training-free and text-based methods have also emerged as computationally efficient alternatives (Schramowski et al., 2023; Yoon et al., 2024; Ban et al., 2024; Armandpour et al., 2023). However, most of these approaches lack a theoretical ground, unlike our work.
@@
Despite these advances, generative models remain susceptible to adversarial prompts, malicious manipulations of learnable parameters, textual cues, or even random noise (Pham et al., 2023; Chin et al., 2024; Zhang et al., 2024b; Tsai et al., 2024). These findings highlight using a single defense such as concept erasing as a standalone solution may be insufficient to ensure safe content generation. We see this as an opportunity for our method to be combined with powerful text-based defense mechanisms to enhance their performance.
@@
The most closely related work is Sparse Repellency (SR) by Kirchhof et al. (2024), a training-free technique that modifies the denoising trajectory to avoid unsafe images $\{\mathbf{x}^{(n)}\}_{n=1}^{N}$. Their denoiser follows $\mathbb{E}_{\text{data}}[\mathbf{x}|\mathbf{x}_{t}]+\sum_{n=1}^{N}\text{ReLU}\left(\tfrac{r}{\|\mathbb{E}_{\text{data}}[\mathbf{x}|\mathbf{x}_{t}]-\mathbf{x}^{(n)}\|}-1\right)\times(\mathbb{E}_{\text{data}}[\mathbf{x}|\mathbf{x}_{t}]-\mathbf{x}^{(n)}).$ The Rectified Linear Unit (ReLU) function ensures that the diffusion trajectory is penalized when the denoiser falls within the neighborhood of radius $r$ around unsafe data, and remains unmodified otherwise. Given a single unsafe image, $\text{ReLU}\left(\tfrac{r}{\|\mathbb{E}_{\text{data}}[\mathbf{x}|\mathbf{x}_{t}]-\mathbf{x}^{(n)}\|}-1\right)(\mathbb{E}_{\text{data}}[\mathbf{x}|\mathbf{x}_{t}]-\mathbf{x}^{(n)})$ resembles the second term in Eq. (4) if the ReLU value is comparable to our $\beta^{*}$. From this, our method can be viewed as a generalization of the SR. However, unlike our method, their guidance does not guarantee sampling from a safe distribution.
@@
Table 1: Comparison of baselines and our method on various datasets. Our method, combined with existing approaches, significantly improves the safety performance while keeping image quality.
@@
(a) Effect on $N$
@@
(c) Effect on prompt weights
Figure 4: Ablation studies of (a) the effect on the number of unsafe data ($N$), (b) the effect on the threshold ($\beta_{t}$), and (c) the effect on the prompt weights. All metrics are evalauted on MMA-Diffusion
Lastly, the work by Biggs et al. (2024) shares a similar theoretical analysis as ours. They propose to merge the weights of DMs separately trained on independent subsets of data, resulting in a sampling distribution that extends beyond the framework outlined in our Theorem. 3.2. However, unlike their method, we do not require additional training of DMs and our analysis defines the safe and unsafe denoisers and their explicit relationship between those.
5 Experiments
In this section, we present the experimental results of our method, Safe Denoiser. Section 5.1 details the outcomes of our text-to-image generation experiments, while the subsequent section explores both unconditional and conditional image generation.
5.1 Text-to-Image Generation
In this section, we conduct an in-depth analysis of the improvements achieved by additionally applying our safe denoiser in text-to-image models. As a baseline, we utilize Stable Diffusion (SD) (Rombach et al., 2022) v1.4222https://huggingface.co/CompVis/stable-diffusion-v1-4. For our experiments, we employ the DDPM sampler.
To assess the model safety, we evaluate Attack Success Rate (ASR) and Toxic Rate (TR) (Yoon et al., 2024). We measure ASR by the proportion of generated images that exceeds 0.6 nude class probability, measured by NudeNet [^3] https://github.com/notAI-tech/NudeNet. The TR is computed by the average of nude class probability, measured also by NudeNet. We select 515 unsafe images as the unsafe dataset of $\{\mathbf{x}^{(1)},\dots,\mathbf{x}^{(N)}\}$ from I2P (Schramowski et al., 2023) that exceeds 0.6 nude class probability.
To evaluate the image quality, we calculate Fréchet Inception Distance (FID) (Heusel et al., 2017) and CLIP (Radford et al., 2021). We use a pytorch package (Seitzer, 2020) to compute the FID by comparing 10K reference images selected from the COCO-2014 (Lin et al., 2014) validation split and 10K generated images from the prompts identically selected from the same COCO dataset. Also, we evaluate the CLIP score using ViT-B-32444https://huggingface.co/openai/clip-vit-base-patch32 with the same dataset.
Quantitative Results
Table  1 summarizes our experimental findings. In these experiments, we utilize unsafe prompts proposed by Ring-A-Bell (Tsai et al., 2024) (79 prompts), UnlearnDiff (Zhang et al., 2024b) (116 sexual prompts), and MMA-Diffusion (Yang et al., 2024) (1000 prompts). For baseline comparisons, we consider both training-based approaches, specifically ESD (Gandikota et al., 2023a) and RECE (Gong et al., 2024), and training-free methods such as SLD (Schramowski et al., 2023) and SAFREE (Yoon et al., 2024). Initially, we observe that using SDv1.4 results in a high percentage of unsafe images across all prompt datasets. As illustrated in Table  1, existing text-based baselines demonstrate performance improvements over SD across each prompt dataset.
Figure 5: Generated samples when negating the Chihuahua class, primarily producing visually similar small dog breeds.
Table 2: Performance evaluation in FFHQ. We use ResNet18 (He et al., 2016a) to classify the sex of generated samples. We compute FID by comparing 1K male subset of CelebA validation and 1K generated images.
Models Female Male FID $\downarrow$

### v4

@@ 3.2 Practial Considerations
For computing Eq. (4), we need to compute three terms: the data denoiser $\mathbb{E}_{\text{data}}[\mathbf{x}|\mathbf{x}_{t}]$, the unsafe denoiser $\mathbb{E}_{\text{unsafe}}[\mathbf{x}|\mathbf{x}_{t}]$ and the weight $\beta^{*}(\mathbf{x}_{t})$. We approximate $\mathbb{E}_{\text{data}}[\mathbf{x}|\mathbf{x}_{t}]$ by utilizing a pre-trained diffusion model. Consequently, the task reduces to deriving $\mathbb{E}_{\text{unsafe}}[\mathbf{x}|\mathbf{x}_{t}]$ and the weight. In this section, we describe our approach to approximating these quantities.
Approximation of the unsafe denoiser.
@@
Each numerator and denominator terms of Eq. (6) approximates the numerator and denominator terms of Eq. (3), respectively. It shows that an unsafe denoiser can be expressed as a weighted sum of the unsafe dataset. Here, the weights $\{\frac{q_{t}(\mathbf{x}_{t}|\mathbf{x}^{(n)})}{\sum_{m=1}^{N}q_{t}(\mathbf{x}_{t}|\mathbf{x}^{(m)})}\}$ form a sum-to-one normalized vector across the unsafe data points, so the unsafe denoiser is approximated as a mixture of unsafe data points.
Approximation of the weight.
Next, we turn our attention to the computation of $\beta^{*}(\mathbf{x}_{t})$ in Eq. (5). Direct calculation is intractable due to the denominator $Z_{\text{safe}}\int p_{\text{safe}}(\mathbf{x})q_{t}(\mathbf{x}_{t}|\mathbf{x})$, which is computationally infeasible [^1] It requies computing $q_{t}(\mathbf{x}_{t}|\mathbf{x})$ over all safe data $\mathbf{x}\sim p_{\text{safe}}(\mathbf{x})$, where safe data includes the entire training dataset excluding few unsafe data. Modern text-to-image models like Stable Diffusion rombach2022high are trained with billions of training data schuhmann2022laion , and is infeasible to iterate the entire data at inference time. to evaluate at every sampling steps. To address this challenge, we approximate $\beta^{*}$ as
@@
$\beta(\mathbf{x}_{t})$ $=\int p_{\text{unsafe}}(\mathbf{x})q_{t}(\mathbf{x}_{t}|\mathbf{x})\mathop{}\!\mathrm{d}\mathbf{x}\approx\frac{1}{N}\sum_{n=1}^{N}q_{t}(\mathbf{x}_{t}|\mathbf{x}^{(n)})$
@@
Input: A pre-trained diffusion model $\bm{\epsilon}_{\bm{\theta}}$; Unsafe data $\{\mathbf{x}^{(n)}\}_{n=1}^{N}$; Hyperparameters $\eta$ and $\beta_{t}$; Critical timesteps $C\subseteq[1,...,T]$; If text-conditional model, positive prompts $\mathbf{c}_{+}$ and unsafe prompts $\tilde{\mathbf{c}}_{-}$
for $t=T$ to 0 do
@@
    Compute $\mathbb{E}_{\text{data}}[\mathbf{x}|\mathbf{x}_{t},\mathbf{c}]$ (e.g., $\mathbf{c}\in\{\tilde{\mathbf{c}}_{+}\}$ for SAFREE or $\mathbf{c}\in\{\mathbf{c}_{+},\tilde{\mathbf{c}}_{-}\}$ for SLD)
@@
    $\mathbf{x}_{0|t}\leftarrow\hat{\mathbb{E}}_{\text{safe}}[\mathbf{x}|\mathbf{x}_{t}]$ (see Eq. (7))
@@
$\begin{split}\hat{\mathbb{E}}_{\text{safe}}[\mathbf{x}|\mathbf{x}_{t}]=\mathbb{E}_{\text{data}}[\mathbf{x}|\mathbf{x}_{t}]+\eta\beta(\mathbf{x}_{t})(\mathbb{E}_{\text{data}}[\mathbf{x}|\mathbf{x}_{t}]-\hat{\mathbb{E}}_{\text{unsafe}}[\mathbf{x}|\mathbf{x}_{t}]),\end{split}$ (7)
@@
3.3 Extending Safe Denoiser to Text-to-Image generation
Table 1: Joint effect of existing text-based guidance (SAFREE) and ours. We evaluate the attack success rate. Both "No" with 0.962 refers to SD-v1.4 rombach2022high with CFG. The lower, the better.
Neg. Prompt
No Yes
Ours No 0.962 0.601
Yes 0.633 0.469
While our methodology is effective as a standalone algorithm, we can also integrate it straightforwardly as a plug-in component into established text-based safety mechanisms, thereby enhancing the overall safety level, as shown in Table  1  [^2] We tested on MMA-Diffusion (yang2024mma, ) nudity prompts and measure the rate the model generates unsafe images. See Section 5 and Table 2 for further details.. For example, when our approach is combined with SAFREE, the predicted clean sample $\mathbf{x}_{0|t}$ (representing the estimated data at step $t=0$ given a sample $\mathbf{x}_{t}$ at step $t$) can be computed by
$\mathbf{x}_{0|t}=\mathbb{E}_{\text{safe}}[\mathbf{x}|\mathbf{x}_{t}]+\underbrace{\lambda(\mathbb{E}_{\text{data}}[\mathbf{x}|\mathbf{x}_{t},\tilde{\mathbf{c}}_{+}]-\mathbb{E}_{\text{data}}[\mathbf{x}|\mathbf{x}_{t}])}_{\text{SAFREE}}.$ (8)
When it is combined with SLD, the formula is as follows:
$\mathbf{x}_{0|t}=\mathbb{E}_{\text{safe}}[\mathbf{x}|\mathbf{x}_{t}]+\underbrace{\lambda(\mathbb{E}_{\text{data}}[\mathbf{x}|\mathbf{x}_{t},\mathbf{c}_{+}]-\mathbb{E}_{\text{data}}[\mathbf{x}|\mathbf{x}_{t}])}_{\text{CFG}}-\underbrace{\mu(\mathbb{E}_{\text{data}}[\mathbf{x}|\mathbf{x}_{t},\tilde{\mathbf{c}}_{-}]-\mathbb{E}_{\text{data}}[\mathbf{x}|\mathbf{x}_{t}])}_{\text{SLD}}.$ (9)
Note these Eq. (8) and Eq. (9) replaces the data denoiser $\mathbb{E}_{\text{data}}[\mathbf{x}|\mathbf{x}_{t}]$ by the safe denoiser $\mathbb{E}_{\text{safe}}[\mathbf{x}|\mathbf{x}_{t}]$, compared to Eq. (1) and Eq. (2), respectively. In implementation, as described in Sec. 3.2, we approximate the safe denoiser by Eq. (7). In diffusion sampling, we utilize this safe $\mathbf{x}_{0|t}$ in either DDPM ho2020denoising or DDIM song2020denoising , see Algorithm 1 for details.
@@ 4 Related Work
Earlier work on machine unlearning in generative modelling focused on object unlearning in classification (forgetting images from a selected class), unconditional image generation (forgetting harmful images) or concept erasing (forgetting harmful concepts). Most of the work belonging to this category required retraining the entire generative models or some part of them, rather than modifying the sampling trajectory or input prompts heng2023selective ; li2024machine ; adapt_then_unlearn ; zhang2024forgetmenot ; gandikota2023unified ; lu2024mace ; gong2024reliable ; lu2024mace . In more recent work, training-free and text-based methods have also emerged as computationally efficient alternatives schramowski2023safe ; yoon2024safree ; ban2024understanding ; armandpour2023reimagine . However, most of these approaches lack a theoretical ground, unlike our work.
@@
Despite these advances, generative models remain susceptible to adversarial prompts, malicious manipulations of learnable parameters, textual cues, or even random noise pham2023circumventing ; chin2024promptingdebugging ; zhang2024generate ; tsai2024ringabell . These findings highlight using a single defense such as concept erasing as a standalone solution may be insufficient to ensure safe content generation. We see this as an opportunity for our method to be combined with powerful text-based defense mechanisms to enhance their performance.
@@
A closely related recent work, Sparse Repellency (SR) kirchhof2024sparse , is a training-free technique that modifies the denoising trajectory to avoid unsafe images. Their denoiser follows $\mathbb{E}_{\text{data}}[\mathbf{x}|\mathbf{x}_{t}]+\sum_{n=1}^{N}\text{ReLU}\left(\tfrac{r}{\|\mathbb{E}_{\text{data}}[\mathbf{x}|\mathbf{x}_{t}]-\mathbf{x}^{(n)}\|}-1\right)\times(\mathbb{E}_{\text{data}}[\mathbf{x}|\mathbf{x}_{t}]-\mathbf{x}^{(n)}).$ ReLU activation ensures that the diffusion trajectory is penalized when the denoiser falls within the neighborhood of radius $r$ around unsafe data, and remains unmodified otherwise. Given a single unsafe image, $\text{ReLU}\left(\tfrac{r}{\|\mathbb{E}_{\text{data}}[\mathbf{x}|\mathbf{x}_{t}]-\mathbf{x}^{(n)}\|}-1\right)(\mathbb{E}_{\text{data}}[\mathbf{x}|\mathbf{x}_{t}]-\mathbf{x}^{(n)})$ resembles the second term, $\mathbb{E}_{\text{safe}}[\mathbf{x}|\mathbf{x}_{t}]$, in Eq. (4) if the ReLU activation is comparable to our $\beta^{*}$. From this point of view, our method can be regarded as a generalization of SR. However, unlike our method, SR does not guarantee that the samples are from a safe distribution.
Diffusion Soup biggs2024diffusion presents a related theoretical analysis by merging DMs trained on separate data subsets, but requires fine-tuning. In contrast, our method is training-free and formally defines safe and unsafe denoisers and their relationship. Dynamic Negative Guidance (DNG) koulischer2025dynamic also uses a similar framework but relies on sequential computation based on a Markov chain and requires extra training for unsafe denoisers, whereas our approach estimates safe and unsafe denoisers in an expectation manner without training overhead.
5 Experiments
We present the experimental results of our method, Safe Denoiser. Section 5.1 details the outcomes of our text-to-image generation experiments, while the subsequent section explores both class-conditional and unconditional image generation.
5.1 Text-to-Image Generation
In this section, we conduct an in-depth analysis of safety issues in text-to-image models, focusing on tasks involving nudity and inappropriate content. The nudity task evaluates how well safety methods prevent harmful outputs as attack difficulty gradually increases. In contrast, the inappropriate content task examines whether these safety methods remain effective when handling multiple concepts simultaneously. We use Stable Diffusion (SD) rombach2022high v1.4333https://huggingface.co/CompVis/stable-diffusion-v1-4 with DDPM sampler. To evaluate safety, we follow previous studies by assessing Attack Success Rate (ASR), Toxic Rate (TR), and Inappropriate Probability (IP) schramowski2023safe ; yoon2024safree . We measure ASR by the proportion of generated images that exceeds 0.6 nude class probability, measured by NudeNet [^4] https://github.com/notAI-tech/NudeNet. The TR is computed by the average of nude class probability, measured also by NudeNet. The IP is the classification probability score of generating inappropriate images, measured by the Q16 classifier (schramowski2022can, ). For the nudity task, we select 515 unsafe images from I2P schramowski2023safe that exceeds 0.6 nude class probability. For the inappropriate content tasks, we randomly sample 3,000 images from I2P as the unsafe dataset. To evaluate, we use the broder dataset CoPro liu2024latent , which covers the same categories of I2P. Notably, all experiments uses the identical unsafe datasets across all baselines for consistency, see Appendix C for details.
Besides safety-related metrics, we prioritize maintaining high image quality and prompt alignment simultaneously. To this end, we calculate Fréchet Inception Distance (FID) heusel2017gans for the generation fidelity and CLIP radford2021learning to measure whether the samples follow human instructions. We use a PyTorch package Seitzer2020FID to compute the FID by comparing 10K reference images selected from the COCO-2014 lin2014microsoft validation split and 10K generated images from the prompts identically selected from the same COCO dataset. Also, we evaluate the CLIP score (radford2021learning, ) using ViT-B-32555https://huggingface.co/openai/clip-vit-base-patch32.
Safe Generation against Nudity Prompts
Table  2 summarizes our experimental findings. In these experiments, we utilize unsafe prompts proposed by Ring-A-Bell tsai2024ringabell (79 prompts), UnlearnDiff zhang2024generate (116 sexual prompts), and MMA-Diffusion yang2024mma (1000 prompts). These prompts are adversarially generated to fool the existing generative models. For baseline comparisons, we consider both training-based approaches, specifically ESD gandikota2023erasing and RECE gong2024reliable , and training-free methods such as SLD schramowski2023safe and SAFREE yoon2024safree . Initially, we observe that about $96.2\%$ of generated SD-v1.4 images are unsafe when using MMA-Diffusion prompts. Existing baselines demonstrate performance improvements over SD across datasets.
Our method, combined with SLD or SAFREE, significantly improves safety performance while maintaining image quality. Notably, the extent of improvement varies considerably depending on the characteristics of the prompts. For instance, with MMA-Diffusion prompts, the performance of text-based baselines (like SLD) is markedly inferior ($88.1\%$ generated images are unsafe) compared to their performance on other prompt datasets such as Ring-A-Bell or UnlearnDiff. This discrepancy arises because MMA-Diffusion prompts lack explicit nudity information due to being part of a white-box adversarial attack, making it challenging for text-based safety methods to erase such concepts. In contrast, our approach employs purely image-based guidance, which results in substantial performance gains from $88.1\%$ to $48.1\%$ in ASR on MMA-Diffusion when combined with existing text-based methods. Our method significantly improves the performance across all other prompt datasets, not limited to MMA-Diffusion.
@@ 5.1 Text-to-Image Generation
Table 2: Performance comparison of baselines on various datasets in safe generation against nudity prompts. Our method, combined with existing approaches, significantly improves the safety performance while keeping image quality.
@@
Table 3: Performance of inappropriate probability (IP) and CLIP Score on the CoPro dataset. Our method incoporating with negative prompts enhances safety performance even across multiple concepts simultaneously.
Method
Harra-
sment $\downarrow$
Hate $\downarrow$
Illegal
Activity $\downarrow$
Self-
harm $\downarrow$
Sexual $\downarrow$
Shock-
ing $\downarrow$
Viole-
nce $\downarrow$
Avg.
IP $\downarrow$
CLIP $\uparrow$
SD-v1.4 0.269 0.154 0.206 0.319 0.120 0.221 0.274 0.223 29.81
+ Ours 0.206 0.148 0.197 0.209 0.109 0.209 0.230 0.187 29.21
SLD 0.223 0.106 0.161 0.247 0.078 0.158 0.217 0.170 29.65
+ Ours 0.168 0.113 0.152 0.169 0.078 0.165 0.212 0.151 28.95
SAFREE 0.182 0.118 0.144 0.183 0.085 0.150 0.206 0.153 28.91
+ Ours 0.156 0.112 0.161 0.153 0.083 0.159 0.185 0.144 28.49
Inappropriate Probability in CoPro Dataset
Table  3 presents that our method consistently achieves enhancement of safe content generation against multiple categories while maintaining a balance in textual prompt alignments across all baselines. Since training-based approaches do not provide official checkpoints for this task, we focus on training-free approaches. Overall, our method effectively improves inappropriate probability (IP) on CoPro dataset. Notably, all baselines show a reduction in average IP when combined with our method. Additionally, our method effectively preserves the alignment between human instructions (prompts) and generated images, with any introduced misalignment being minimal, as demonstrated by CLIP scores. Furthurmore, our method performs on-par with previous methods in terms of the sample-wise aesthetic scores, showing that there is a minimal impact in the sample quality by applying our method, see Appendix E. These results suggest that our method effectively manages multiple concepts simultaneously while reliably generations away from unsafe content.
Ablation Studies
We present a pair of ablation studies to evaluate the robustness and effectiveness of our method. First, Figure 4(a) shows the effect of the number of unsafe data points on model performance. We observe that increasing the number of unsafe data points leads to better performance.
(a) Effect on NN
@@
Figure 4: Ablation studies of (a) the effect on the number of unsafe data (NN), (b) the effect on the threshold ($\beta_{t}$).
Table 4: Experiments of the class negation on ImageNet. Top-1* is the classification accuracy of the generated samples on the negated class (Chihuahua). Refer Appendix D and E.
Method Prec $\uparrow$ Rec $\uparrow$ Top-1* $\downarrow$
Baseline 0.72 0.63 0.68
B + SR 0.59 0.54 0.00
B + Ours 0.62 0.58 0.00
We then explore the influence of the threshold parameter $\beta_{t}$ (see Algorithm 1), which governs the application of the safe denoiser. For simplicity, we fixed $\beta_{t}$ across all time steps. Figure 4(b) shows the performance exhibits a U-shaped relationship to $\beta_{t}$. Specifically, when $\beta_{t}=0$, the safe denoiser is applied to all samples $\mathbf{x}_{t}$ regardless of their safety status. Conversely, when $\beta_{t}=\infty$, the safe denoiser is not applied. At intermediate values of $\beta_{t}$, the safe denoiser is applied selectively to a certain proportion of unsafe samples $\mathbf{x}_{t}$. The U-shaped trend indicates that selectively applying the safe denoiser to unsafe samples based on an appropriate $\beta_{t}$ value is optimal, thereby balancing denoising efficacy and computational efficiency. Additional ablation studies are presented in Appendix to discuss in-depth analysis of the scalability, robustness, and effectiveness of our methods.
Computation Overhead
Table 5: Wall-clock time.
Models
Time
(s/img)
SD-v1.4 3.18
+ Ours ($N=515$) 3.20
SAFREE 4.22
+ Ours ($N=515$) 4.24
+ Ours ($N=3,000$) 4.29
Table  5 presents the wall-clock time for image generation on NVIDIA RTX4090 with 24GB memory. Thanks to GPU parallelism, the additional time introduced by our method scales sub-linearly since modern GPUs optimize batched matrix multiplications with efficient job scheduling. For example, our method increases from 4.22s to 4.29s when using $3,000$ negative images (an overhead of only 0.07s), while increasing from 4.22s to 4.24s when using $515$ images. Conversely, SAFREE adds over 1s per image. Given that SAFREE and ours perform similarly in Table  1, our method shows a better performance-efficiency curve.
5.2 Class-Conditional and Unconditional Generation with Safe Denoiser
Table 6: Performance in FFHQ. We use ResNet18 (he2016deep, ) to classify the sex of generated samples. We compute FID by comparing male data and generated images.
Models Female $\downarrow$ Male $\uparrow$ FID $\downarrow$